OpenStudy (juscallmesteve):

I could use some help to solve this problem. Lim X^(tanx) X-->0 I am fairly sure I would need to use l'hospitals rule but not sure how to convert the x^tanx into a fraction

OpenStudy (phi):

yes, this one is tricky. one way is write x = \( e^{\ln x}\)

OpenStudy (phi):

so your problem is a limit for \[ x^{\tan x} = \left( e^{\ln x}\right)^{\tan x}= e^{\ln x \cdot \tan x} \]

OpenStudy (juscallmesteve):

is there a rule that will show me why you did what you did?

OpenStudy (phi):

now you do \[ \lim_{x \rightarrow 0} \ln x \tan x \]

OpenStudy (reemii):

Other way of presenting phi's trick: \(x^{\tan x} = e^{\ln(x^{\tan x})} = e^{\tan x \ln x}\). Same result.

OpenStudy (phi):

by definition \[ e^{\ln x }= x \]

OpenStudy (juscallmesteve):

Lol I honestly did not know that

OpenStudy (phi):

meanwhile, the expression is now a product, so we still have to massage it into a quotient

OpenStudy (juscallmesteve):

So I take apply l'hospitals to lim lnxtanx X-->0

OpenStudy (juscallmesteve):

So Lnx/tanx^-1 ?

OpenStudy (phi):

I haven't worked it out. It will take some fiddling to get this.

OpenStudy (juscallmesteve):

Ok, quick question What can I search in Google to teach me how to solve this? Converting x^tanx into a fraction

OpenStudy (juscallmesteve):

Or is there a rule I need to know? Because product rule wouldn't work.

OpenStudy (phi):

you can only try. I would not give up (yet) on what we are doing

OpenStudy (juscallmesteve):

Ok will return

OpenStudy (reemii):

Did you apply L'Hospital rule already?

OpenStudy (juscallmesteve):

Lol I did this wrong entirely I applied the product rule to the exponent I got e^(lnxsec^2x + tanx/x)

OpenStudy (juscallmesteve):

This one is very confusing for me, I will need to find some YouTube videos

OpenStudy (reemii):

on \(\frac{\ln x}{\frac{\cos x}{\sin x}}\), I obtain, \[ \lim_{x\to 0} \frac{1/x}{-1/\sin^2x} = \lim_{x\to0} -\frac{\sin^2x}{x} \] If you know that \(\frac{\sin x}{x}\to 1\) as \(x\to 0\), then you're done. Otherwise, apply L'Hospital rule again.

OpenStudy (phi):

yes, that looks good.

OpenStudy (phi):

so the original limit is e^0 = 1

OpenStudy (juscallmesteve):

So essentially what you are doing is changing x^tanx To lnX/tanx And tanx = cosx/sinx ?

OpenStudy (juscallmesteve):

And then going from there LnX = 1/x

OpenStudy (juscallmesteve):

Is that an identity going from Cosx/sinx to -1/sin^2x

OpenStudy (phi):

that is the derivative

OpenStudy (juscallmesteve):

I lol That just didn't click that we went from lnX to 1/x which was done by taking the derivative. So it would make sense to take derivative of the bottom

OpenStudy (reemii):

\(\ln x\tan x = \frac{\ln x}{\frac{\cos x}{\sin x}}\)

OpenStudy (phi):

if we summarize, we get to ln x tan x = ln x / cot x this is now in the form oo/oo take the derivatives: 1/x and -1/sin^x and rewrite that as - sin^2(x )/ x if you don't the the sin x / x limit, then apply again: -2 sin x cos x / 1 and the limit is -2*0*1= 0 and that 0 is the limit of \[ e^{\lim \ln x \tan x} = e^0 = 1\]

OpenStudy (juscallmesteve):

Ok thank you, once closed can I come back to this question?

OpenStudy (juscallmesteve):

Like to look over?

OpenStudy (reemii):

I think it's possible. Also, you can find it easily on your profile page.